# Searching for quantum-gravity footprint around stellar-mass black holes

### v1

@@ 4.1 Cygnus X-1
This is the oldest and best known stellar-mass black hole. It has the most reliable measured values of $M=21.1_{-2.3}^{+2.2}M_{\odot}$, $d=2.22_{-0.17}^{+0.18}$ kpc, $i=27.51_{-0.57}^{+0.77}$ deg, and $a=0.9696-0.9985$ [41]. It is then possible to calculate $r_{\rm g}=31\pm 3$ km and $r_{\rm isco}=(1.21-1.74)r_{\rm g}$. The innermost radius of the accretion disk in units of [$r_{\rm g}$] for different observations is displayed in Fig. 2.
Figure 2: Cygnus X-1: inner radius of the accretion disk in units of [$r_{\rm g}$]. The two dashed grey lines represent the range of $r_{\rm isco}$ as expected from GR. Reference number refers to the source of data: 1: [42]; 2: [43]; 3-4: [44]; 5: [45]; 6: [46]; 7: [47]; 8-11: [48]; 12-13 [49]; 14: [50].
The smallest value of $r_{\rm in}=(1.29_{-0.27}^{+0.28})r_{\rm g}$ – still consistent with the expectations from GR – was derived from a simultaneous observation with Suzaku and NuSTAR ($1-300$ keV) between October 31 and November 1, 2012 [46]. We considered the best fit model 4 in Table 2, where the normalization of diskbb is reported to be $N=20800_{-800}^{+1200}$, from which we can estimate $r_{\rm in}$ via Eq. (2.7). Unless explicitly stated otherwise, we always assumed $\kappa=1.7$.
@@ 4.2 GRS 1915+105
This is one of the most monitored and studied stellar-mass black hole. Also in this case, there are reliable measured values of $M=11.8\pm 0.6M_{\odot}$, $d=9.4\pm 1.0$ kpc, $i=64\pm 4$ deg, and $a=0.970-0.997$ [51, 52, 53]. It is then possible to calculate $r_{\rm g}=17.4\pm 0.9$ km and $r_{\rm isco}=(1.28-1.74)r_{\rm g}$. The innermost radius of the accretion disk in units of [$r_{\rm g}$] for different observations is displayed in Fig. 3.
Figure 3: GRS 1915+105: inner radius of the accretion disk in units of [$r_{\rm g}$]. The two dashed grey lines represent the range of $r_{\rm isco}$ as expected from GR. Reference number refers to the source of data: 1-2: [54]; 3: [55]; 4-7: [56]; 8-10: [57]; 11: [58]; 12: [59]; 13-22: [60]; 23-27: [61]; 28-29: [62]; 30-32: [63]; 33-36: [64]; 37-46: [65]; 47-50: [66]; 51-54: [67]; 55-60: [68]; 61: [69]; 62: [70]; 63-68: [71]; 69: [72].
This is the only case where we adopted $\kappa=1.9$, as suggested by [65, 69], since the color temperature is generally higher than usual for a stellar-mass black hole ($T_{\rm col}\sim 1-2$ keV).
Also in this case, $r_{\rm in}$ is consistent with the expectations from general relativity within the measurement errors. Some cases are borderline: one reason can be that the diskbb does not include relativistic corrections (gravitational redshift and Doppler boosting), which can become significant for high inclinations [19]. We remind that $i=64\pm 4$ deg for GRS 1915+105 [53].
@@ 4.3 XTE J1550-564
This black hole is interesting, because it is not maximally rotating. Its spin has been estimated to be $a=0.49_{-0.20}^{+0.13}$ [73]. The other quantities necessary to estimate the reference and measured values of the inner radii are $M=9.10\pm 0.61M_{\odot}$, $d=4.38_{-0.41}^{+0.58}$ kpc, and $i=74.7\pm 3.8$ deg [74]. The gravitational radius is $r_{\rm g}=13.4\pm 0.9$ km, and $r_{\rm isco}=(3.74-5.01)r_{\rm g}$. The innermost radius of the accretion disk in units of [$r_{\rm g}$] for different observations is displayed in Fig. 4.
@@
Figure 4: XTE J1550-564: inner radius of the accretion disk in units of [$r_{\rm g}$]. The two dashed grey lines represent the range of $r_{\rm isco}$ as expected from GR. Reference number refers to the source of data: 1: [75]; 2-13: [76]; 14-15: [77]; 16: [78]; 17: [79]; 18-20: [80]; 21: [81]; 22: [82]; 23-27: [83].
@@
In this case, there are more significant deviations (reference numbers 14, 15, 18, 21). The cases 14 and 15 were from [77] and refer to RXTE observations in the $3-200$ keV energy band between April and June 2000. Table 1 of the cited paper reported the apparent radius calculated from the normalization of diskbb by assuming $d=6$ kpc and $i=73.1$ deg. The fits were done every one-two days and a thermal component is reported in 21 over 43 fits, with $T_{\rm col}\sim 0.37-0.95$ keV. We can already note that such temperature can be biased by the $3$ keV low-energy threshold of the detector. Since we are interested in searching for deviations from the expectations of general relativity, we selected the two smallest apparent radii, which are $r_{\rm col}\sim 32.5$ km (MJD $51662.2$), and $r_{\rm col}\sim 39.4$ km (MJD $51673.4$). There is also a third value of $r_{\rm col}\sim 39$ km at MJD $51682.3$, but the photon index of the power-law model is harder than those of the previous two cases ($1.76$ vs $2.33$, and $2.31$). This suggests the presence of a significant Comptonization, which can alter the estimate of the apparent radius [28], and therefore we did not consider it. For the two remaining cases, it is necessary to understand if they were measured during a high/soft state, with negligible corona contribution. As already written, the photon indexes of the power-law models were steep, but there are no indication of the fluxes due to the different components. We found some indications only in Fig. 4 in [77], which plots the fluxes in the $2-50$ keV energy band of the two models (power law vs diskbb). One point corresponding to MJD $51662.2$ is explicitly indicated: in this case, the power-law flux is about $2-3$ times that from the accretion disk. The other point, corresponding to MJD $51673.4$, is not indicated. However, by looking at the other points in the plot, one can reasonably think that the power-law component should have been at least comparable to the thermal emission. Therefore, we can conclude that these two inner radii estimates do not refer to high/soft states, and are affected by a strong Comptonization, which led to an underestimation of the value [28]. Similar cases are the number 18 from [80] (it is the period 7 in Table 1), and the number 21 from [81] (it is the case C, model 2 in Table 2).
@@
A last note of caveat: recently, the viewing angle has been challenged by [82]. They suggested a value of $i\sim 40$ deg, much smaller than previously measured. This would have a significant impact on the estimation of the inner radius. The value calculated for $i\sim 75$ deg is $r_{\rm in}\sim 5.90r_{\rm g}$ (reference number 22), but in the case of $i\sim 40$ deg, then $r_{\rm in}\sim 3.46r_{\rm g}$, smaller than expected from general relativity ($r_{\rm isco,min}\sim 3.74r_{\rm g}$). This new viewing angle seems quite unlikely, because the kinematic analysis of the relativistic jet sets a tight constrain on the viewing angle, indicating a value of $i\sim 71^{\circ}$ [84].
Figure 5: GX 339-4: inner radius of the accretion disk in units of [$r_{\rm g}$]. The two dashed grey lines represent the range of $r_{\rm isco}$ as expected from GR. Reference number refers to the source of data: 1-2: [86]; 3: [87]; 4-5: [88]; 6: [89]; 7: [90]; 8-10: [91]; 11-13: [92]; 14-18: [93]; 19: [94]; 20: [95]; 21: [96]; 22: [97]; 23: [98]; 24-25: [99]; 26: [100]; 27-29: [101]; 30-33: [102]; 34-36: [103]; 38-40: [105]; 41: [106]; 42: [107]; 43: [108].
@@ 4.4 GX 339-4
Yet another well-known and observed stellar-mass black hole. Its reference quantities are: $M=9.0_{-1.2}^{+1.6}M_{\odot}$, $d=8.4\pm 0.9$ kpc, $i=30\pm 1$ deg, and $a=0.95_{-0.08}^{+0.02}$ [85]. The gravitational radius is $r_{\rm g}=13\pm 2$ km, and $r_{\rm isco}=(1.73-2.51)r_{\rm g}$. The innermost radius of the accretion disk in units of [$r_{\rm g}$] for different observations is displayed in Fig. 5.
Also in this case, there are no significant deviations. The only value outside the expected range – but stil consistent within the measurement errors – is the number 14 [93]. It refers to a INTEGRAL and XMM-Newton observation ($0.7-200$ keV) performed between January and March 2007. The value of $r_{\rm in}=1.21_{-0.29}^{+0.98}r_{\rm g}$ has been estimated from the data of Epoch 1 of Table 2. We noted the presence also of a hard component modelled with a power law with $\Gamma\sim 1.46$, and Table 3 stated that the disk component accounts for $\sim 21$% of the total flux. Again, a significant Comptonization led to an underestimate of the inner radius of the disk.
@@ 4.5 XTE J1650-500
As we wrote in the Introduction, we expect that the most significant deviations from GR may occur around the smallest black holes. Therefore, we decided to study also two objects with the smallest mass, although with large uncertainties and the lack of some reference information. The first one is XTE J165050016505001650-5001650 - 500 and has its reference quantities are: $M=4.0\pm 0.6M_{\odot}$, $d=2.6\pm 0.7$ kpc, $i=70\pm 4$ deg, and $a\sim 0.9982$ [109, 110, 111]. The gravitational radius is $r_{\rm g}=5.9\pm 0.9$ km, and $r_{\rm isco}\sim 1.23r_{\rm g}$.
We found only two estimates in the literature, both showing no deviations from the expected value of $r_{\rm isco}$: $r_{\rm in}=(18\pm 8)r_{\rm g}$ [112], and $r_{\rm in}=(5.3\pm 1.7)r_{\rm g}$ [113].
Searching for some unpublished data in the archives, we found one Swfit observation (id 00031762001) performed on July 17, 2010. We analysed the data by using HEASoft v. 6.33.2 software package and CALDB updated on August 28, 2024. Swift/XRT data were reduced by using standard procedures (xrtpipeline with default values). We found no sources, with a flux upper limit of $\sim 3\times 10^{-11}$ erg cm-2 s-1 ($3\sigma$ detection, $\sim 1.0$ ks exposure in photon counting mode, $N_{\rm H}=3.68\times 10^{21}$ cm-2 [114], power-law model with $\Gamma=2$). This means that XTE J1650-500 was likely in a quiescent state.
@@ 4.6 GRO J0422+32
The last object is the candidate to be the smallest black hole known, with $M=2.7_{-0.5}^{+0.7}M_{\odot}$, $d=2.49\pm 0.30$ kpc, $i=55.6\pm 4.1$ deg [115, 116]. No information about the spin was found, and therefore it is not possible to estimate the expected $r_{\rm isco}$. We found only one paper useful to estimate the inner radius: $r_{\rm in}=(5.1\pm 2.3)r_{\rm g}$ [117]. We also found one Swift observation (id 00032976001) performed on October 3, 2013. We analysed the data with the same procedure adopted for the above cited case, but found no source, with a a flux upper limit of $\sim 1.5\times 10^{-11}$ erg cm-2 s-1 ($3\sigma$ detection, $\sim 1.2$ ks exposure in photon counting mode, $N_{\rm H}=1.51\times 10^{21}$ cm-2 [114], power-law model with $\Gamma=2$). Also GRO J0422+32 was likely in a quiescent state.

### v2

@@ 4.1 Cygnus X-1
This is the oldest and best known stellar-mass black hole. It has the most reliable measured values of $M=21.1_{-2.3}^{+2.2}M_{\odot}$, $d=2.22_{-0.17}^{+0.18}$ kpc, $i=27.51_{-0.57}^{+0.77}$ deg, and $a=0.9696-0.9985$ (Miller-Jones et al., 2021). It is then possible to calculate $r_{\rm g}=31\pm 3$ km and $r_{\rm isco}=(1.21-1.74)r_{\rm g}$. The inner radius of the accretion disk in units of [$r_{\rm g}$] for different observations is displayed in Fig. 2.
Figure 2: Cygnus X-1: inner radius of the accretion disk in units of [$r_{\rm g}$]. The two dashed grey lines represent the range of $r_{\rm isco}$ as expected from GR, while the continuous line represents the $r_{\rm isco}$ with the maximum allowable spin of $a=0.99$ according to Mummery (2025). Reference number refers to the source of data: 1: Dotani et al. (1997); 2: Poutanen et al. (1997); 3-4: Cui et al. (1998); 5: Frontera et al. (2001); 6: Tomsick et al. (2014); 7: Sugimoto et al. (2016); 8-11: Walton et al. (2016); 12-13 Kushwaha et al. (2021); 14: Yan et al. (2021).
The smallest value of $r_{\rm in}=(1.29_{-0.27}^{+0.28})r_{\rm g}$ – still consistent with the expectations from GR – was derived from a simultaneous observation with Suzaku and NuSTAR ($1-300$ keV) between October 31 and November 1, 2012, by Tomsick et al. (2014). We considered the best fit model 4 in Table 2, where the normalization of diskbb is reported to be $N=20800_{-800}^{+1200}$, from which we can estimate $r_{\rm in}$ via Eq. (7). Unless explicitly stated otherwise, we always assumed $\kappa=1.7$.
Kushwaha et al. (2021) reported spectral fits with both diskbb and kerrbb. The latter has fixed the $M$, $i$, and $d$ to values similar to those by Miller-Jones et al. (2021) and obtained a lower limit to the spin $a>0.9878$ for the Observation G ($r_{\rm isco}<1.49r_{\rm g}$), which is our number 12 in Fig. 2, corresponding to $r_{\rm in}=(1.66\pm 0.36)r_{\rm g}$. Kushwaha et al. (2021) reported also another measurement for the Observation I ($a>0.9843$, $r_{\rm isco}<1.55r_{\rm g}$), which we did not include in Fig. 2 as diskbb gives a values greater than Obs G ($r_{\rm in}=1.89_{-0.41}^{+0.40}r_{\rm g}$). The two values are consistent with each other and with the expected value from the Kerr metric. Other authors (Kawano et al., 2017; Zdziarski et al., 2024a, b) adopted the kerrbb model to estimate the spin from high/soft state observations, when $r_{\rm in}\sim r_{\rm isco}$, and they obtain values consistent with those of Miller-Jones et al. (2021).
@@ 4.2 GRS 1915+105
This is one of the most monitored and studied stellar-mass black hole. Also in this case, there are reliable measured values of $M=11.8\pm 0.6M_{\odot}$, $d=9.4\pm 1.0$ kpc, $i=64\pm 4$ deg, and $a=0.970-0.997$ (Miller et al., 2013; Sreehari et al., 2020; Reid & Miller-Jones, 2023). It is then possible to calculate $r_{\rm g}=17.4\pm 0.9$ km and $r_{\rm isco}=(1.28-1.74)r_{\rm g}$. The inner radius of the accretion disk in units of [$r_{\rm g}$] for different observations is displayed in Fig. 3.
Figure 3: GRS 1915+105: inner radius of the accretion disk in units of [$r_{\rm g}$]. The two dashed grey lines represent the range of $r_{\rm isco}$ as expected from GR, while the continuous line represents the $r_{\rm isco}$ with the maximum allowable spin of $a=0.99$ according to Mummery (2025). The arrows indicate upper limits. Reference number refers to the source of data: 1-2: Taam et al. (1997); 3: Muno et al. (1999); 4-7: Feroci et al. (1999); 8-10: Rao et al. (2000); 11: Belloni et al. (2000); 12: Zdziarski et al. (2001); 13-22: Vadawale et al. (2001); 23-27: Ueda et al. (2002); 28-29: Naik et al. (2002); 30-32: Done et al. (2004); 33-36: Ohkawa & Kitamoto (2005); 37-46: Rodriguez et al. (2008); 47-50: Vierdayanti et al. (2010); 51-54: Ueda et al. (2010); 55-60: Rahoui et al. (2010); 61: Neilsen et al. (2011); 62: Miller et al. (2016); 63-68: Mineo et al. (2017); 69: HESS Collaboration (2018).
This is the only case where we adopted $\kappa=1.9$, as suggested by Rodriguez et al. (2008) and Neilsen et al. (2011), since the color temperature is generally higher than usual for a stellar-mass black hole ($T_{\rm col}\sim 1-2$ keV). The values reported by Rodriguez et al. (2008) – indicated in Fig. 3 by the numbers from 37 to 46 – are interesting, because they adopted the ezdiskbb model. It was not possible to recalculate the normalization, because the inclination value is missing; therefore, we calculate $r_{\rm in}$ by using the peak temperature and the flux of the accretion disk from their Table 1. Obviously, in this case, we did not apply the $\varsigma$ correction factor by Kubota et al. (1998), because the ezdiskbb model includes the zero-torque boundary condition.
Also for GRS 1915+105, $r_{\rm in}$ is consistent with the expectations from general relativity within the measurement errors. Some cases are borderline: one reason can be that the diskbb does not include relativistic corrections (gravitational redshift and Doppler boosting), which can become significant for high inclinations (Ebisawa et al., 2003). We remind that $i=64\pm 4$ deg for GRS 1915+105 (Reid & Miller-Jones, 2023).
@@ 4.3 XTE J1550-564
This black hole is interesting, because it is not maximally rotating. Its spin has been estimated to be $a=0.49_{-0.20}^{+0.13}$ (Steiner et al., 2011). The other quantities necessary to estimate the reference and measured values of the inner radii are $M=9.10\pm 0.61M_{\odot}$, $d=4.38_{-0.41}^{+0.58}$ kpc, and $i=74.7\pm 3.8$ deg (Orosz et al., 2011). The gravitational radius is $r_{\rm g}=13.4\pm 0.9$ km, and $r_{\rm isco}=(3.74-5.01)r_{\rm g}$. The inner radius of the accretion disk in units of [$r_{\rm g}$] for different observations is displayed in Fig. 4.
@@
Figure 4: XTE J1550-564: inner radius of the accretion disk in units of [$r_{\rm g}$]. The two dashed grey lines represent the range of $r_{\rm isco}$ as expected from GR, while the continuous line represents the $r_{\rm isco}$ with the maximum allowable spin of $a=0.99$ according to Mummery (2025). Reference number refers to the source of data: 1: Sobczak et al. (1999); 2-13: Sobczak et al. (2000); 14-15: Rodriguez et al. (2003); 16: Miller et al. (2003); 17: Kubota & Done (2004); 18-20: Kubota & Makishima (2004); 21: Sriram et al. (2016); 22: Connors et al. (2019); 23-27: Connors et al. (2020).
@@
In this case, there are more significant deviations (reference numbers 14, 15, 18, 21). We should have removed these observations, but we thought they can be a good example of how the presence of Comptonization can alter the measurement of $r_{\rm in}$. The cases 14 and 15 were from Rodriguez et al. (2003) and refer to RXTE observations in the $3-200$ keV energy band between April and June 2000. Table 1 of the cited paper reported the apparent radius calculated from the normalization of diskbb by assuming $d=6$ kpc and $i=73.1$ deg. The fits were done every one-two days and a thermal component is reported in 21 over 43 fits, with $T_{\rm col}\sim 0.37-0.95$ keV. We can already note that such temperature can be biased by the $3$ keV low-energy threshold of the detector. Since we are interested in searching for deviations from the expectations of general relativity, we selected the two smallest apparent radii, which are $r_{\rm col}\sim 32.5$ km (MJD $51662.2$), and $r_{\rm col}\sim 39.4$ km (MJD $51673.4$). There is also a third value of $r_{\rm col}\sim 39$ km at MJD $51682.3$, but the photon index of the power-law model is harder than those of the previous two cases ($1.76$ vs $2.33$, and $2.31$). This suggests the presence of a significant Comptonization, which can alter the estimate of the apparent radius (Merloni et al., 2000), and therefore we did not consider it. For the two remaining cases, it is necessary to understand if they were measured during a high/soft state, with negligible corona contribution. As already written, the photon indexes of the power-law models were steep, but there are no indication of the fluxes due to the different components. We found some indications only in Fig. 4 in Rodriguez et al. (2003), which plots the fluxes in the $2-50$ keV energy band of the two models (power law vs diskbb). One point corresponding to MJD $51662.2$ is explicitly indicated: in this case, the power-law flux is about $2-3$ times that from the accretion disk. The other point, corresponding to MJD $51673.4$, is not indicated. However, by looking at the other points in the plot, one can reasonably think that the power-law component should have been at least comparable to the thermal emission. Therefore, we can conclude that these two inner radii estimates do not refer to high/soft states, and are affected by a strong Comptonization, which led to an underestimation of the value (Merloni et al., 2000). Similar cases are the number 18 from Kubota & Makishima (2004) (it is the period 7 in Table 1), and the number 21 from Sriram et al. (2016) (it is the case C, model 2 in Table 2).
@@
A final note of warning: recently, the viewing angle has been challenged by Connors et al. (2019). They suggested a value of $i\sim 40$ deg, much smaller than previously measured. This would have a significant impact on the estimation of the inner radius. The value calculated for $i\sim 75$ deg is $r_{\rm in}\sim 5.90r_{\rm g}$ (reference number 22), but in the case of $i\sim 40$ deg, then $r_{\rm in}\sim 3.46r_{\rm g}$, smaller than expected from general relativity ($r_{\rm isco,min}\sim 3.74r_{\rm g}$). This new angle seems rather unlikely, because the kinematic analysis of the relativistic jet imposes a tight constraint on the viewing angle, indicating a value of $i\sim 71^{\circ}$ (Steiner & McClintock, 2012).
Figure 5: GX 339-4: inner radius of the accretion disk in units of [$r_{\rm g}$]. The two dashed grey lines represent the range of $r_{\rm isco}$ as expected from GR, while the continuous line represents the $r_{\rm isco}$ with the maximum allowable spin of $a=0.99$ according to Mummery (2025). Reference number refers to the source of data: 1-2: Miller et al. (2004a); 3: Miller et al. (2004b); 4-5: Belloni et al. (2006); 6: Reis et al. (2008); 7: Miller et al. (2008); 8-10: Del Santo et al. (2008); 11-13: Motta et al. (2009); 14-18: Caballero-García et al. (2009); 19: Shidatsu et al. (2011); 20: Motta et al. (2011); 21: Tamura et al. (2012); 22: Rahoui et al. (2012); 23: Plant et al. (2014); 24-25: Ludlam et al. (2015); 26: Kubota & Done (2016); 27-29: Stiele & Kong (2017); 30-33: Sridhar et al. (2020); 34-36: Shui et al. (2021); 38-40: Yang et al. (2023); 41: Peirano et al. (2023); 42: Liu et al. (2023); 43: Jana et al. (2024).
@@ 4.4 GX 339-4
Yet another well-known and observed stellar-mass black hole. Its reference quantities are: $M=9.0_{-1.2}^{+1.6}M_{\odot}$, $d=8.4\pm 0.9$ kpc, $i=30\pm 1$ deg, and $a=0.95_{-0.08}^{+0.02}$ Parker et al. (2016). The gravitational radius is $r_{\rm g}=13\pm 2$ km, and $r_{\rm isco}=(1.73-2.51)r_{\rm g}$. The inner radius of the accretion disk in units of [$r_{\rm g}$] for different observations is displayed in Fig. 5.
Also in this case, there are no significant deviations. The only value outside the expected range – but stil consistent within the measurement errors – is the number 14 (Caballero-García et al., 2009). It refers to a INTEGRAL and XMM-Newton observation ($0.7-200$ keV) performed between January and March 2007. The value of $r_{\rm in}=1.21_{-0.29}^{+0.98}r_{\rm g}$ has been estimated from the data of Epoch 1 of Table 2. We noted the presence also of a hard component modelled with a power law with $\Gamma\sim 1.46$, and Table 3 stated that the disk component accounts for $\sim 21$% of the total flux. Again, a significant Comptonization led to an underestimate of the inner radius of the disk.
@@ 4.5 XTE J1650-500
As we wrote in the Introduction, we expect the most significant deviations from general relativity to occur around the smallest black holes. Therefore, we decided to also study two objects with the smallest masses, although with large uncertainties and the lack of some reference information. The first one is XTE J165050016505001650-5001650 - 500 and has its reference quantities are: $M=4.0\pm 0.6M_{\odot}$, $d=2.6\pm 0.7$ kpc, $i=70\pm 4$ deg, and $a\sim 0.9982$ (Orosz et al., 2004; Homan et al., 2006; Slaný & Stuchlík, 2008). The gravitational radius is $r_{\rm g}=5.9\pm 0.9$ km, and $r_{\rm isco}\sim 1.23r_{\rm g}$.
We found only two estimates in the literature, both showing no deviations from the expected value of $r_{\rm isco}$: $r_{\rm in}=(18\pm 8)r_{\rm g}$ Miller et al. (2002), and $r_{\rm in}=(5.3\pm 1.7)r_{\rm g}$ (Miniutti et al., 2004).
Searching for some unpublished data in the archives, we found one Swfit observation (id 00031762001) performed on July 17, 2010. We analysed the data by using HEASoft v. 6.33.2 software package and CALDB updated on August 28, 2024. Swift/XRT data were reduced by using standard procedures (xrtpipeline with default values). We found no sources, with a flux upper limit of $\sim 3\times 10^{-11}$ erg cm-2 s-1 ($3\sigma$ detection, $\sim 1.0$ ks exposure in photon counting mode, $N_{\rm H}=3.68\times 10^{21}$ cm-2 HI4PI Collaboration (2016), power-law model with $\Gamma=2$). This means that XTE J1650-500 was likely in a quiescent state.
@@ 4.6 GRO J0422+32
The latest object is a candidate for the smallest black hole known, with $M=2.7_{-0.5}^{+0.7}M_{\odot}$, $d=2.49\pm 0.30$ kpc, $i=55.6\pm 4.1$ deg (Gelino & Harrison, 2003; Casares et al., 2022). No information about the spin was found, and therefore it is not possible to estimate the expected $r_{\rm isco}$. We found only one paper useful to estimate the inner radius: $r_{\rm in}=(5.1\pm 2.3)r_{\rm g}$ (Shrader et al., 1997). We also found one Swift observation (id 00032976001) performed on October 3, 2013. We analyzed the data with the same procedure adopted for the above case, but we did not find any source, with an upper limit of the flux of $\sim 1.5\times 10^{-11}$ erg cm-2 s-1 ($3\sigma$ detection, $\sim 1.2$ ks exposure in photon counting mode, $N_{\rm H}=1.51\times 10^{21}$ cm-2 HI4PI Collaboration (2016), power-law model with $\Gamma=2$). Also GRO J0422+32 was likely in a quiescent state.